# SpatialCrafter: Unleashing the Imagination of Video Diffusion Models for Scene Reconstruction from Limited Observations

### v1

@@ Abstract
Novel view synthesis (NVS) boosts immersive experiences in computer vision and graphics. Existing techniques, though progressed, rely on dense multi-view observations, restricting their application. This work takes on the challenge of reconstructing photorealistic 3D scenes from sparse or single-view inputs. We introduce SpatialCrafter, a framework that leverages the rich knowledge in video diffusion models to generate plausible additional observations, thereby alleviating reconstruction ambiguity. Through a trainable camera encoder and an epipolar attention mechanism for explicit geometric constraints, we achieve precise camera control and 3D consistency, further reinforced by a unified scale estimation strategy to handle scale discrepancies across datasets. Furthermore, by integrating monocular depth priors with semantic features in the video latent space, our framework directly regresses 3D Gaussian primitives and efficiently processes long-sequence features using a hybrid network structure. Extensive experiments show our method enhances sparse view reconstruction and restores the realistic appearance of 3D scenes.
@@ 1 Introduction
The complexity of this problem stems from three main aspects: first, the diversity and complexity of real-world scene make precise modeling extremely challenging; second, occluded regions in sparse viewpoints are difficult to recover; and finally, limited geometric cues constrain the accuracy of reconstruction. Compared to traditional methods that depend on numerous viewpoints, achieving high-quality reconstruction from sparse views requires stronger prior knowledge and more efficient reasoning capabilities, which is the core challenge of this problem. To address these challenges, some methods [29, 81, 26, 39, 18, 77] constrain the optimization process by introducing regularization terms, but these methods need to be optimized for each scene, are not generalizable, and are difficult to cope with complex scenarios. Recently, some feed-forward methods [8, 9, 56, 48, 22, 53, 73, 74] have achieved generalizable sparse view scene reconstruction by directly predicting 3D Gaussian primitives for each pixel. However, such methods produce severe artifacts and implausible images when reconstructing occluded regions and applying them to the extrapolation setting. To address this challenge, some methods [34, 60, 43, 61] have introduced the priors from the diffusion model to achieve sparse or single-view NVS, but they perform poorly on scene-level data and lack precise pose control and 3D consistency. This is due to the fact that low-dimensional camera information such as Euler angles and extrinsic struggle to provide comprehensive control signals to the generative models.
In this paper, we propose a framework named SpatialCrafter for scene reconstruction and novel view synthesis based on sparse or single-view inputs. Our innovation lies in leveraging the rich physical world knowledge embedded in video diffusion models to provide plausible additional observations for scene reconstruction, thereby effectively reducing the problem complexity. Specifically, we first focus on enhancing precise camera control and 3D consistency in generated videos. To achieve this, we parameterize camera settings using ray embeddings or metric depth-warped images and incorporate them into the video diffusion model via a trainable camera encoder. Furthermore, we propose an epipolar attention mechanism that improves 3D consistency between video frames through explicit geometric constraints. Previous methods [60, 20] often experience performance degradation and difficulty generating large-motion videos when trained on multiple scene datasets. We identified that these issues primarily stem from scale ambiguity in scene datasets. To overcome this, we introduce a unified scale estimator to calibrate scene dataset, enabling effective joint training across multiple datasets.
@@ 2 Related Work
Figure 2: Overview of our pipeline. Our generative reconstruction pipeline consists of two parts: camera-controlled video generation, and then reconstruction. First, we set the exploration path based on the input views. In the video generation module, ray embedding is used to parameterize the camera information, and a cross-attention mechanism with epipolar constraints is introduced to improve the 3D consistency of the generated video. The robust 3D scene reconstruction pipeline integrates monocular depth priors with semantic features extracted from the video latent space, and directly regress 3D Gaussian primitives through a feed-forward manner.
@@ 2.1 Sparse-View Scene Reconstruction
NeRF [58] and Gaussian Splatting  [25] have achieved photorealistic representations of 3D scenes, but require optimization on densely collected image sets for each scene. To address this problem, some approaches [29, 18, 81, 67] focus on constraining the optimization process of 3D representations via manually designed regularization. Some other methods [23, 8, 9, 52, 65, 68] are trained on large-scale datasets and directly predict the 3D representation in a feed-forward manner. Recently, some approaches [14, 33, 48, 30] based on end-to-end 3D reconstruction models [56, 28], have achieved efficient and high-quality 3D reconstruction. However, these methods cannot handle occluded areas, and when reconstruction fails (e.g., in the case of large viewpoint changes), the results of subsequent novel view synthesis can be significantly compromised.
@@ 2.3 Controllable Video Generation
While recent text-to-video diffusion models have achieved remarkable progress, they inherit the controllability limitations of their text-to-image counterparts, often requiring additional conditioning mechanisms to align generated videos with user intent. Several works [57, 19, 24, 66] have been carried out to introduce a variety of conditions into video generation models. Recently, attention has focused on controlling the camera motion of the video. Some approaches [19, 51] employs motion-specific training for predefined camera movements, though this framework struggles with complex trajectory synthesis. Some methods [60, 20, 63] enable more complex camera control by parameterizing the camera trajectory and injecting it into a pre-trained video diffusion model via a trainable camera encoder. MotionCtrl [60] controls camera and object motions independently through learnable adapters, while CameraCtrl [20] introduces Plücker embeddings for accurate 6-DoF camera pose representation. AC3D [1] optimizes the pose adjustment scheme to accelerate convergence and limits the injection of camera features to specific locations. ViewCrafter [71] and ReconX [33] pushes this direction by enabling pixel-accurate view compositing through point-based rendering formulation.
@@ 3 Method
This section begins by explaining our camera-controlled video generation model (Section 3.1), which employs ray embeddings for precise pose control. Then, we describe how to use epipolar geometry constraints to enhance 3D consistency in the video frames. Next, we introduce ourscene reconstruction pipeline (Section 3.2), which aims to perform stable 3D reconstruction based on the generated video latents. These innovations collectively improve the accuracy and consistency of 3D reconstructions from generated video.
@@ 3.1 Camera-Conditioned Video Generation
Scale Alignment. While recent methods [60, 20] have achieved promising results in novel view synthesis, their reliance on ground-truth camera extrinsic from Re10K [79] during inference inherently limits their applicability to in-the-wild scenarios. Moreover, the metric-scaled poses in Re10K create alignment challenges when jointly training with other datasets (e.g., DL3DV [31]) where poses are estimated using COLMAP, primarily due to scale ambiguity issues. To address this issue, we introduced a metric depth estimation model [5] to achieve scale alignment of camera poses between different datasets. During the data preprocessing stage, we align the metric 3D point cloud predicted by the model with the sparse point cloud generated using COLMAP to calculate the scale factor, thereby obtaining absolute-scale camera extrinsics.
Camera Injection. Inspired by recent methods [20, 72, 47], we chose to use ray embeddings [20] or the depth warping frames [71] to represent the camera information. This over-parameterized approach [72] provides more accurate pose information to the video generator than other methods. Specifically, for a ray defined by the origin $o\in\mathbb{R}^{3}$ and the normalized direction $d\in\mathbb{R}^{3}$, we can represent it as $(o\times d,d)\in\mathbb{R}^{6}$. Given the camera parameters, the direction of the ray $d^{\prime}$ corresponding to the pixel coordinates $(u,v)$ can be calculated as $d^{\prime}=\mathbf{R}\mathbf{K}^{-1}(u,v,1)^{\top}+\mathbf{T}$. For depth warping frames, we leverage the pretrained depth estimation model [55] to map the reference image into a 3D point cloud. Subsequently, we render novel views at specified camera poses by projecting this point cloud via the target camera parameters. Afterwards, we use a trainable conditional encoder to inject the camera information into the video model instead of fine-tuning all model parameters. The training objective is:
$\mathcal{L}=\mathbb{E}_{z,z_{0},\epsilon,\boldsymbol{P},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},t,\phi({c}))\right\|_{2}^{2}\right]$ (1)
where $\boldsymbol{P}$ represents the ray embeddings, $z_{0}$ denotes the latent embeddings of the reference frame, and $t$ indicates the timestamps.
@@
$\mathcal{L}=\mathbb{E}_{z,z_{0},z_{n}\epsilon,\boldsymbol{P},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},z_{n},t,\boldsymbol{P})\right\|_{2}^{2}\right]$ (2)
where $z_{n}$ and $z_{0}$ are the latent embeddings of the final and initial frames, respectively. During the inference phase, we use DUSt3R to estimate the extrinsic and intrinsic parameters for the input views.
@@ 3.2 Video-based Scene Reconstruction
Table 1: Quantitative comparison to the camera conditioned video generation method on RealEstate10K, DL3DV, and Tanks-and-Temples dataset.
Latent Feature Fusion. Given input video latents $\mathbf{z}\in\mathbb{R}^{T\times H\times W\times C}$ and camera poses encoded as ray embeddings $\mathbf{p}\in\mathbb{R}^{T\times H\times W\times 6}$, we transform them into token sequences through a patchification process. Specifically, we apply spatial patchification to latent features to obtain latent tokens $\mathbf{z}_{t}\in\mathbb{R}^{N\times d_{z}}$, while ray embeddings undergo 3D-patchification to produce pose tokens $\mathbf{p}_{t}\in\mathbb{R}^{N\times d_{p}}$. To incorporate explicit geometric guidance, we leverage monocular depth estimation within our pipeline. For each input RGB video frame, we estimate dense depth maps $\mathbf{D}\in\mathbb{R}^{T\times H\times W\times 1}$, which are processed by a dedicated depth encoder to yield depth tokens $\mathbf{d}_{t}\in\mathbb{R}^{N\times d_{d}}$. The three token sets are channel-wise concatenated to form a unified representation $\mathbf{x}=[\mathbf{z}_{t};\mathbf{p}_{t};\mathbf{d}_{t}]\in\mathbb{R}^{N\times(d_{z}+d_{p}+d_{d})}$, which is then linearly projected to a lower-dimensional space $\mathbf{x}^{\prime}\in\mathbb{R}^{N\times d}$ before being fed into a sequence of mamba and transformer blocks. Mamba [17] achieves the same token sequence processing functionality as Transformer, but reduces computational complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L)$, making it particularly suitable for dense reconstruction tasks. In our Mamba blocks, we implement bi-directional scanning across the token sequence. First, we compute state parameters $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{R}^{N\times d}$ from the input tokens using a linear projection. Then, we execute the State Space Model (SSM) in both forward ($\mathbf{y}_{f}$) and backward ($\mathbf{y}_{b}$) directions. The final output of each Mamba block is computed as $\mathbf{y}=\mathbf{y}_{f}+\mathbf{y}_{b}$ followed by a final linear transformation.
@@ 4.2.1 Controllable Video Generation
Table 4: Quantitative comparison on the RealEstate10K dataset. Small and Large refer to input images with low and high overlap ratios, respectively. Greater overlap means greater temporal coherence between adjacent images.
4.2.2 Sparse View Reconstruction
@@ 4.2.3 Single View 3D Generation
Comparison. As shown in Table 5, our method consistently achieves superior performance on image quality metrics compared to the baselines. For a fair comparison with ZeroNVS [43] and LucidDreamer [10], which are limited to square image inputs, we crop the generated novel views before computing the quantitative metrics. Figure 6 shows that scenes reconstructed using our method have less noise and distortion in occluded areas.
@@ 4.3 Ablation Study
Figure 9: Qualitative comparison of video results. Our method outperforms other video generation models by reducing the shot transition rate and ensuring smooth transitions, particularly when faced with large differences in input viewing angles.
4.4 Discussions
Existing commercial video generation models like Kling and GEN-3 Turbo face a critical challenge: significant differences in input viewing angles often result in unnatural camera transitions. As shown in Figure 9, our approach effectively addresses this issue by fine-tuning the video generation model on extensive 3D scene datasets and integrating specific 3D information during the inference process. This enables our model to produce videos with more continuous and natural camera transitions.
@@ 5 Conclusion
We present SpatialCrafter, a framework for scene reconstruction and novel view synthesis from sparse or single-view inputs. By leveraging video diffusion models to generate plausible additional observations, we effectively reduce the complexity of sparse-view scene reconstruction. Our key contributions include: (1) precise camera control via ray embeddings or depth-warped images with a trainable condition encoder; (2) a unified scale estimator solving the scale ambiguity in multi-dataset training; and (3) a hybrid Mamba-Transformer architecture that combines monocular depth priors with semantic features from video latent space to directly regress 3D Gaussian primitives. Experiments show our method outperforms existing methods, especially in single-view extrapolation and scenarios with little overlaps. Future work will focus on handling extreme complex geometries, and extending to dynamic scene reconstruction.
Appendix A Implementation Details
A.1 Training Details
Our video generation model is built upon SVD [3], an image-video diffusion model based on UNet. The training process incorporates a continuous-time noise scheduler (Karras et al., 2022), enabling the model to learn to gradually denoise high-variance Gaussian noise. Specifically, we define the noisy data as $x_{t}=x_{0}+n(t)$, where the added Gaussian noise $n(t)\sim\mathcal{N}(0,\sigma^{2}(t)I)$. The iterative denoising process corresponds to the probability flow ODE:
$dx=-\dot{\sigma}(t)\sigma(t)\nabla_{x}\log p(x;\sigma(t))dt,$
where $\nabla_{x}\log p(x;\sigma(t))$ is the score function parameterized by a neural network $D_{\theta}$. The training objective focuses on denoising score matching: $E\left[\|D_{\theta}(x_{0}+n;\sigma,c)-x_{0}\|^{2}_{2}\right]$, with $c$ representing the conditioning information. We follow the EDM framework, parameterizing the learnable denoiser as $D_{\theta}(x;\sigma)=c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(c_{\text{in}}(\sigma)x;c_{\text{noise}}(\sigma))$.
In our work, we use a relative camera system to convert all camera poses relative to the first frame located at the world origin. First, we train the model at a resolution of $320\times 512$ with a frame length of 25, for 50,000 iterations, and we perform center cropping on the input images to standardize the resolution. Subsequently, we adjust the model to a higher resolution of $576\times 1024$ and continue training for another 10,000 iterations, setting the learning rate to $51e-$ 5 and using the Adam optimizer with a 1,000-step warm-up. The training is conducted on 8 NVIDIA A800 GPUs with a batch size of 16. It is important to note that we only train the camera encoder, epipolar attention modules, and temporal attention while keeping the weights of other parts frozen. We first train on the Re10K and ACID datasets, with the sampling interval of video frames linearly increasing from 2 to 10. After that, we incorporate the DL3DV dataset, randomly sampling the video frame intervals between $[2,10]$ to enable the model to adapt to different motion speeds. We choose Lightning as the training framework, utilizing mixed precision fp16 and DeepSpeed ZeRO-2 to enhance efficiency. During inference, we implement the DDIM sampler [49] alongside classifier-free guidance [21] to enhance performance.
A.2 Evaluation Details
Video Generation. When using COLMAP to extract the camera pose of the generated video, the reliability of the obtained pose is often low. Therefore, we employ DUSt3R for a more robust pose estimation. To compare with previous methods, we transform the camera coordinates of the estimated poses to be relative to the first frame and normalize the scale using the furthest frame. We then calculate the rotation distance relative to the ground truth rotation matrices of each generated novel view sequence, expressed as $R_{\text{dist}}=\sum_{i=1}^{n}\arccos\left(\frac{\text{tr}(R_{\text{gen}}^{i}R_{\text{gt}}^{i^{T}})-1}{2}\right)$. We also compute the $T_{dist}=\sum_{i=1}^{n}\|T_{gt}^{i}-T_{gen}^{i}\|_{2}$, which measures the Euclidean distance between ground truth and estimated translations across $n$ video frames.
Sparse View Reconstruction. To ensure a fair comparison with feed-forward baselines [8, 9], we conducted experiments on $256\times 256$ resolutions. To comprehensively evaluate our method, we compared the results of input views under small and large overlap ratios. We employed RoMa to determine the overlap ratio of image pairs. First, we obtained dense matches from the first image to the second and vice versa, considering pixels with matching scores exceeding 0.005 as valid. Next, we calculated the overlap ratio for each image pair by dividing the number of valid matched pixels by the total number of pixels, with the final overlap ratio defined as the minimum of the two directional calculations. Additionally, to compare with optimization-based reconstruction baselines [14, 29], we used only 2 ground truth training images per scene and evaluated using 12 views. The evaluation metrics included PSNR, SSIM, and LPIPS.
Appendix B Discussions
B.1 Comparison with Very Recent Methods
ViewCrafter. ViewCrafter [71] leverages point cloud rendering as a condition for fine-tuning video diffusion models. However, constrained by Dust3R’s capabilities, the approach may encounter challenges in synthesizing novel views with extensive view range transformations, particularly when generating front-view images from limited back-view inputs. Moreover, in scenarios where DUSt3R fails—such as incorrectly predicting planar geometries—ViewCrafter’s performance can be significantly compromised. In contrast, our proposed generation module decouples from the reconstruction module, effectively mitigating these limitations and enhancing view synthesis robustness.
DimensionX. DimensionX [51] achieves promising results by fine-tuning respective LORA models for different motion patterns. However, this method has some limitations: first, it cannot generate more complex motion patterns; second, it lacks effective control over the speed and amplitude of motion. In contrast, our method introduces camera embedding as the conditional signal, which effectively solves the above problems, thereby enabling the generation of more complex motion patterns and providing precise control over the speed and amplitude of motion.
ReconX. ReconX [33] initially constructs a global point cloud representation and encodes it into a contextual space, serving as 3D geometric conditional information for video diffusion models. However, the method’s performance critically depends on the geometric reconstruction accuracy of the underlying DUST3R algorithm. When DUST3R fails to precisely align and reconstruct geometric structures, the point cloud information injected into the video diffusion model may introduce significant geometric distortions, leading to suboptimal generation results. More importantly, the current approach is limited to multi-view interpolation scenarios, lacking an effective solution for single-view extrapolation tasks
CAT3D. CAT3D [16] extends the text-to-image (T2I) generation model by a large-scale data-driven method to learn a 3D prior of the scene, which usually requires a large amount of training time. In addition, the generation stage requires rendering a large number of frames (about 100 frames) to reconstruct the scene. We recognize that video generation models trained on large video datasets embed inherent scene prior information, and therefore propose a new method based on fine-tuning with a video diffusion model. Our method speeds up convergence and retains the generative prior of the pre-trained video model, thereby enhancing the model’s generalization ability. By introducing explicit geometric constraints, we further improve the 3D consistency of the generated scenes. Notably, during the inference stage, we utilize the geometric prior for initialization and reconstruction, greatly reducing the number of images required and the optimization time.
Appendix C More Qualitative Results
C.1 More Generated Video Results
As shown in Figure 14, we present a series of generated video sequences covering different scenarios, including indoor environments, outdoor landscapes, and object-centered scenes. Our method shows excellent stability in generating video frame sequences with high 3D consistency, and can effectively predict the plausible appearance details of previously unseen regions. In addition, Figure 12 demonstrates the view interpolation ability of our video generation model using sparse input views, and the performance remains stable even when the input view angles differ greatly.
C.2 Basic Camera Trajectories
Our method demonstrates promising generalization ability in camera-controllable video generation, especially for common motion patterns such as zoom-in, zoom-out, and panning. As shown in Figure 10, our method can generate stable videos with basic camera trajectories using a large-scale training dataset that contains a diverse and complex distribution of camera motion patterns. In addition, in Figure 11, we also show the control ability of our method for different motion amplitudes.
C.3 Single View Exploration Results
Figure 13 illustrates that scene reconstruction from sparse input views often suffers from severe occlusions and geometric gaps due to limited observation angles. By implementing a multi-view observation strategy using video diffusion models, we can progressively fill in these missing regions, thereby enhancing the reconstruction’s completeness and geometric detail. As we continuously increase and synthesize observation perspectives, the scene’s geometric structure becomes more complete, significantly improving the quality and accuracy of the 3D reconstruction. In Figure 11, we show that by controlling the camera parameters, we can directly control the magnitude of the movement, which is general in different scenarios.
C.4 More 3D Reconstruction Results
Figure 16 presents additional reconstruction and rendering results based on generated video frames, demonstrating that our method effectively achieves high-quality reconstructions in indoor, outdoor, and object-centered scenes.
Appendix D Limitations
Our method encounters several challenges in video generation and scene reconstruction. First, existing video diffusion models suffer from significant inference efficiency bottlenecks, particularly when processing large-scale scene reconstruction tasks requiring numerous video frames, leading to prohibitively high computational costs. Second, current video diffusion models are constrained in the number of generated video frames, which impedes comprehensive scene reconstruction for complex scenarios. Lastly, the lack of effective user interaction mechanisms represents a substantial limitation, as users cannot exert fine-grained control over the generation process and results, restricting the model’s adaptability in practical applications.
Figure 10: Visualization of our method for generating videos with basic motion trajectories.
Figure 11: Visualization of our method for controlling different motion amplitudes.
Figure 12: More Visualization of interpolated video based on sparse view input.
Figure 13: Visualization of exploration and reconstruction based on single-view input.
Figure 14: More video generation results with high 3D consistency based on single-view input.
Figure 15: Consistent frames generated by our approach. Our method demonstrates strong generalization capability across diverse scenarios, including the DL3DV [31], Tanks-and-Temples benchmark [27], and synthetic images from generative models, while maintaining geometric coherence and structural fidelity.
Figure 16: More video generation and reconstruction results with high 3D consistency based on single-view input.

### v2

@@ Abstract
Novel view synthesis (NVS) boosts immersive experiences in computer vision and graphics. Existing techniques, though progressed, rely on dense multi-view observations, restricting their application. We tackle the task of reconstructing photorealistic 3D scenes from only one or a few input views. We introduce SpatialCrafter, a framework that leverages the rich knowledge in video diffusion models to generate plausible additional observations, thereby alleviating reconstruction ambiguity. Through a trainable camera encoder and an epipolar attention mechanism for explicit geometric constraints, we achieve precise camera control and 3D consistency, further reinforced by a unified scale estimation strategy to handle scale discrepancies across datasets. Furthermore, by integrating monocular depth priors with semantic features in the video latent space, our framework directly regresses 3D Gaussian primitives and efficiently processes long-sequence features using a hybrid network structure. Extensive experiments show our method enhances sparse view reconstruction and restores the realistic appearance of 3D scenes. Project page: https://franklinz233.github.io/projects/spatialcrafter/.
@@ 1 Introduction
This problem is hard because (1) real-world scenes are diverse and complex, (2) occluded regions are unseen in sparse captures, and (3) geometric cues are too few to constrain reconstruction. To address these challenges, some methods [29, 88, 26, 40, 17, 84] constrain the optimization process by introducing regularization terms, but these methods need to be optimized for each scene, are not generalizable, and are difficult to cope with complex scenarios. Recently, some feed-forward methods [7, 8, 60, 50, 22, 55, 79, 80] have achieved generalizable sparse view scene reconstruction by directly predicting 3D Gaussian primitives for each pixel. However, such methods produce severe artifacts when reconstructing occluded regions and applying them to the extrapolation setting. To address this challenge, some methods [35, 64, 45, 65] have introduced the priors from the diffusion model to achieve sparse or single-view NVS, but they perform poorly on scene-level data and lack precise pose control and 3D consistency. This is due to the fact that low-dimensional camera information such as Euler angles and extrinsic struggle to provide comprehensive control signals to the generative models.
In this paper, we propose a framework named SpatialCrafter for scene reconstruction and novel view synthesis based on sparse or single-view inputs. Our innovation lies in leveraging the rich physical world knowledge embedded in video diffusion models to provide plausible additional observations for scene reconstruction, thereby effectively reducing the problem complexity. Specifically, we first focus on enhancing precise camera control and 3D consistency in generated videos. To achieve this, we parameterize camera settings using ray embeddings [78] or metric depth-warped images [77] and incorporate them into the video diffusion model via a trainable camera encoder. Furthermore, we propose an epipolar attention mechanism that improves 3D consistency between video frames through explicit geometric constraints. Previous methods [64, 19] often experience performance degradation and difficulty generating large-motion videos when trained on multiple scene datasets. We identified that these issues primarily stem from scale ambiguity in scene datasets. To overcome this, we introduce a unified scale estimator to calibrate scene dataset, enabling effective joint training across multiple datasets.
@@ 2 Related Work
Figure 2: Overview of our pipeline. Our generative reconstruction pipeline consists of two parts: camera-controlled video generation, and then video-based reconstruction. First, we set the exploration path based on the input views. In the video generation module, ray embeddings are used to parameterize the camera trajectories, and a cross-attention mechanism with epipolar constraints is introduced to improve the 3D consistency of the generated video. The video-based scene reconstruction pipeline integrates monocular depth priors with semantic features extracted from the video latent space, and directly regress 3D Gaussian primitives via a feed-forward manner.
@@ 2.1 Sparse-View Scene Reconstruction
NeRF [62] and Gaussian Splatting  [25] have achieved photorealistic representations of 3D scenes, but require optimization on densely collected image sets for each scene. To address this problem, some approaches [29, 17, 88, 71] focus on constraining the optimization process of 3D representations via manually designed regularization. Some other methods [23, 7, 8, 54, 69, 72, 58, 83] are trained on large-scale datasets and directly predict the 3D representation in a feed-forward manner. Recently, some approaches [13, 50] based on end-to-end 3D reconstruction models [60, 28, 7], have achieved efficient and high-quality 3D reconstruction. However, these methods cannot handle occluded regions, and reconstruction failures (e.g., under large viewpoint changes) can severely affect the quality of novel view synthesis. More relevant to our work, some works [57, 34, 31] have employed priors from video diffusion models to enhance the performance of sparse-view reconstruction.
@@ 2.3 Controllable Video Generation
While recent text-to-video diffusion models have achieved remarkable progress, they inherit the controllability limitations of their text-to-image counterparts, often requiring additional conditioning mechanisms to align generated videos with user intent. Several works [61, 18, 24, 70] have been carried out to introduce a variety of conditions into video generation models. Recently, attention has focused on controlling the camera motion of the video. Some approaches [18, 53] employs motion-specific training for predefined camera movements, though this framework struggles with complex trajectory synthesis. Some methods [64, 19, 20, 67, 1] enable more complex camera control by parameterizing the camera trajectory and injecting it into a pre-trained video diffusion model via a trainable camera encoder. Another line of work [77, 34, 74, 43] advances this direction by enabling pixel-accurate view compositing through a point-based rendering approach.
@@ 3 Method
This section begins by explaining our camera-controlled video generation model (Section 3.1), which employs ray embeddings for precise pose control. Then, we describe how to use epipolar geometry constraints to enhance 3D consistency in the video frames. Next, we introduce ourscene reconstruction pipeline (Section 3.2), which aims to perform stable 3D reconstruction based on the generated video latents.
@@ 3.1 Camera-Conditioned Video Generation
Scale Alignment. To address this challenge, our approach first employs VGGT [58] to estimate initial camera parameters, represented by the extrinsic matrix $\mathbf{P}=[\mathbf{R}|\mathbf{T}]$, and a corresponding depth map $d_{v}$ for each video frame. To establish a consistent scale across these datasets, we leverage the Metric3D [73] to infer a canonical metric depth map, denoted $d_{m}$. We then align our initial depth prediction, $d_{v}$, to the canonical scale by calculating a scale factor $s$, defined as the ratio of the inter-percentile ranges (IPR) of the two depth maps: $s=\operatorname{IPR}_{0.8,0.2}(d_{v})/\operatorname{IPR}_{0.8,0.2}(d_{m})$. This scale factor is then used to adjust the translation component of the camera extrinsics, resulting in $\mathbf{P}_{scaled}=[\mathbf{R}|s\cdot\mathbf{T}]$. This procedure yields camera extrinsics with a consistent absolute scale, effectively resolving alignment issues when fusing data from different sources.
Camera Injection. Inspired by recent methods [19, 78, 49], we chose to use ray embeddings [19] or the depth warping frames [77] to represent the camera information. Specifically, for a ray defined by the origin $o\in\mathbb{R}^{3}$ and the normalized direction $d\in\mathbb{R}^{3}$, we can represent it as $(o\times d,d)\in\mathbb{R}^{6}$. Given the camera parameters, the direction of the ray $d^{\prime}$ corresponding to the pixel coordinates $(u,v)$ can be calculated as $d^{\prime}=\mathbf{R}\mathbf{K}^{-1}(u,v,1)^{\top}+\mathbf{T}$. For depth warping frames, we leverage the pretrained depth estimation model [59] to map the reference image into a 3D point cloud. Subsequently, we render novel views at specified camera poses by projecting this point cloud via the target camera parameters. Afterwards, we use a trainable conditional encoder to inject the camera information into the video model instead of fine-tuning all model parameters. The training objective is:
$\mathcal{L}=\mathbb{E}_{z,z_{0},\epsilon,{C},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},t,\phi({C}))\right\|_{2}^{2}\right]$ (1)
where $\phi({C})$ represents the camera conditional encodings, $z_{0}$ denotes the latent embeddings of the reference frame, and $t$ indicates the timestamps.
@@
$\mathcal{L}=\mathbb{E}_{z,z_{0},z_{n}\epsilon,{C},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},z_{n},t,\phi({C})))\right\|_{2}^{2}\right]$ (2)
where $z_{n}$ and $z_{0}$ are the latent embeddings of the final and initial frames, respectively. During the inference phase, we use VGGT [58] to estimate the extrinsic and intrinsic parameters for the input views.
@@ 3.2 Video-based Scene Reconstruction
Latent Feature Fusion. Given input video latents $\mathbf{z}\in\mathbb{R}^{T\times H\times W\times C}$ and camera poses encoded as ray embeddings $\mathbf{p}\in\mathbb{R}^{T\times H\times W\times 6}$, we transform them into token sequences through a patchification process. Specifically, we apply spatial patchification to latent features to obtain latent tokens $\mathbf{z}_{t}\in\mathbb{R}^{N\times d_{z}}$, while ray embeddings undergo 3D-patchification to produce pose tokens $\mathbf{p}_{t}\in\mathbb{R}^{N\times d_{p}}$. To incorporate explicit geometric guidance, we leverage monocular depth estimation within our pipeline. For each input RGB video frame, we estimate dense depth maps $\mathbf{D}\in\mathbb{R}^{T\times H\times W\times 1}$, which are processed by a dedicated depth encoder to yield depth tokens $\mathbf{d}_{t}\in\mathbb{R}^{N\times d_{d}}$. The three token sets are channel-wise concatenated to form a unified representation $\mathbf{x}=[\mathbf{z}_{t};\mathbf{p}_{t};\mathbf{d}_{t}]\in\mathbb{R}^{N\times(d_{z}+d_{p}+d_{d})}$, which is then linearly projected to a lower-dimensional space $\mathbf{x}^{\prime}\in\mathbb{R}^{N\times d}$ before being fed into a sequence of mamba and transformer blocks. Mamba [16] achieves the same token sequence processing functionality as Transformer, but reduces computational complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L)$, making it particularly suitable for dense reconstruction tasks. In our Mamba blocks, we implement bi-directional scanning across the token sequence. First, we compute state parameters $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{R}^{N\times d}$ from the input tokens using a linear projection. Then, we execute the State Space Model in both forward ($\mathbf{y}_{f}$) and backward ($\mathbf{y}_{b}$) directions. The final output of each Mamba block is computed as $\mathbf{y}=\mathbf{y}_{f}+\mathbf{y}_{b}$ followed by a final linear transformation.
@@
Table 1: Quantitative comparison to the camera conditioned video generation method on RealEstate10K [86], DL3DV [32], and Tanks-and-Temples [27] dataset.
@@ 4.2.1 Controllable Video Generation
Table 4: Quantitative comparison on the RealEstate10K [86] dataset. Small and Large refer to input images with low and high overlap ratios, respectively. Greater overlap means greater temporal coherence between adjacent images.
@@
4.2.2 Sparse View Reconstruction
Figure 7: Additional visualizations of generated videos and their corresponding 3DGS representations.
@@ 4.2.3 Single View 3D Generation
Comparison. As shown in Table 5, our method consistently achieves superior performance on image quality metrics compared to the baselines. For a fair comparison with ZeroNVS [45] and LucidDreamer [9], which are limited to square image inputs, we crop the generated novel views before computing the quantitative metrics. Figure 6 shows that scenes reconstructed using our method have less noise and distortion in occluded areas. For more visualizations, see Figure 7.
@@ 5 Conclusion
We present SpatialCrafter, a framework for scene reconstruction and novel view synthesis from sparse or single-view inputs. By leveraging video diffusion models to generate plausible additional observations, we effectively reduce the complexity of sparse-view scene reconstruction. Our key contributions include: (1) precise camera control via ray embeddings or depth-warped images with a trainable condition encoder; (2) a unified scale estimator solving the scale ambiguity in multi-dataset training; and (3) a hybrid Mamba-Transformer architecture that combines monocular depth priors with semantic features from video latent space to directly regress 3D Gaussian primitives. Experiments show our method outperforms existing methods, especially in single-view extrapolation and scenarios with little overlaps. Future work will focus on extending to dynamic scene.